OpenStudy (solomonzelman):

Just trying out... (the math there is just for me see the entirety of my latex)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \left.\int_{a}^{b} f(x)={\rm F}(x)\right|^{b}_{a}={\rm F}(b)-{\rm F}(a) }\)

OpenStudy (solomonzelman):

Successful! ``` \(\large\color{black}{ \displaystyle \left.\int_{a}^{b} f(x) ={\rm F}(x)\right|^{b}_{a}={\rm F}(b)-{\rm F}(a) }\) ``` \(\color{red}{\text{\left.}}\) integral, blah blah \(\color{red}{\text{\right|}}\) \(\color{black}{\Uparrow}\) (Placed \(\color{red}{\text{\left.}}\) before the integral, and that works to make the line higher to make it look nicer)

OpenStudy (liv1234):

@SolomonZelman Can you teach me how to do this if you have some time?

OpenStudy (solomonzelman):

I am not that good :)

OpenStudy (liv1234):

It seems hard to do haha, so you seemed to do really good

OpenStudy (solomonzelman):

\(\large\color{#A44999}{ \displaystyle \oint F(x){~\sf d}{\bf \ell}}\) Nice colors, http://www.w3schools.com/tags/ref_colorpicker.asp choose or make up. Again I am just trying out things on here

OpenStudy (solomonzelman):

\(\large{\bbox[4px,#fff1ff ,border:2px solid red ]{\color{#660033}{ \left.\begin{matrix} ~~~x &~~ & i^x~ \\ \\[0.9em] ~~4k & | & i^{4k}=(i^4)^x=1^x=\color{blue}{1} \\[0.9em] ~~4k+1 & | & i^{4k+1}=i^{4k}\cdot i^{1} =1\cdot i =\color{blue}{i}\\[0.9em] ~~4k+2& | & i^{4k+2}=i^{4k} \cdot i^{2}=1\cdot(-1)=\color{blue}{-1}\\[0.9em] ~~ 4k+3 & | & i^{4k+3}=i^{4k} \cdot i^{3}=i^{4k} \cdot i^{2}\times i^{9} \\[0.9em]&&=1\cdot(-1)\cdot i =\color{blue}{-i} \\[0.9em] \end{matrix}\right. }}}\)

OpenStudy (solomonzelman):

\(\large\color{#003366}{\displaystyle\lim_{x \rightarrow ~\infty }\left[\frac{x}{\ln(x!)}\right] }\) \(\large\color{#003366}{\displaystyle\lim_{x \rightarrow ~\infty }\left[\frac{x\ln(e)}{\ln(x!)}\right] }\) \(\large\color{#003366}{\displaystyle\lim_{x \rightarrow ~\infty }\left[\frac{\ln(e^x)}{\ln(x!)}\right] }\) \(\large\color{#003366}{\displaystyle \ln \left[\lim_{x \rightarrow ~\infty }\frac{e^x}{x!}\right]=0 }\)

OpenStudy (solomonzelman):

\(\large\color{#660033}{\displaystyle\lim_{x \rightarrow ~\infty }\left[\frac{\left( \ln x\right)^{k+2}}{x}\right]\quad ;\quad k\in \mathbb{N} }\) \(\large\color{#660033}{\displaystyle\lim_{x \rightarrow ~\infty }\left[\frac{~\frac{d}{dx}~\left[\left( \ln x\right)^{k+2} \right]~}{\frac{dx}{dx}}\right] }\) \(\large\color{#660033}{\displaystyle\lim_{x \rightarrow ~\infty }\left[\frac{~(k+2)\left( \ln x\right)^{k+1}~\times \dfrac{1}{x} }{1}\right] }\) \(\large\color{#660033}{\displaystyle\lim_{x \rightarrow ~\infty }\left[\frac{~(k+2)\left( \ln x\right)^{k+1} }{x}\right]}\) and differentiate a (k+2) number of times... and this is how it goes. \(\large\color{#339966}{\displaystyle\lim_{x \rightarrow ~\infty }\left[\frac{~\left( \ln x\right)^{k+2} }{x}\right]}\) \(\large\color{#339966}{\displaystyle\lim_{x \rightarrow ~\infty }\left[\frac{~\dfrac{d^{k+2}}{dx^{k+2}}~\left( \ln x\right)^{k+2} }{1}\right]}\) \(\large\color{#339966}{\displaystyle\lim_{x \rightarrow ~\infty }\left[\frac{(k+2)! }{x}\right]=0}\)

OpenStudy (solomonzelman):

the colors are awesome... (squeeze trm to prove this for non integer k, suppose that there is a positive noninteger S, and a positive integer A that is greater than S, such that: k<k+S<k+A and since when k+A is in the power of ln x, and when k is in the power of ln x, the limit is 0, therefore when any nuninteger positive number is in the power of ln, the limit is 0)

OpenStudy (solomonzelman):

just in case, for math completion.